100$ is put into an account that earns 3.5% interest compunded annually. How much would be in the account after 10 years? Also how long would it take for the account to have $200? For the first part I got 145$ and the second part I had no idea!thank you so much!
Answer by stanbon(75887)

100$ is put into an account that earns 3.5% interest compunded annually. How much would be in the account after 10 years?
A(10) = 100(1+(0.035/1))^(1*10) = 100(1.035)^10 = $141.06
=============================
Also how long would it take for the account to have $200?
Solve for "t":
200 = 100(1.035)^t
-----
2 = 1.035^t
t = log(2)/log(1.035)
t = 20.15 years
--------------------
Cheers,
Stan H.